If a vehicle travels a fixed distance at 30 miles per hour in 10 minutes, how fast must it travel to cover the same distance in 15 minutes?
Answer by ewatrrr(24785)


Hi,
Let x represent speed to cover same distance in 15 minutes
30mph*(1/6)hr = x*(1/4)hr
5 = x/4
20mph = x